### v1

# Impact of LLMs news Sentiment Analysis on Stock Price Movement Prediction

Abstract
This paper addresses stock price movement prediction by leveraging LLM-based news sentiment analysis. Earlier works have largely focused on proposing and assessing sentiment analysis models and stock movement prediction methods, however, separately. Although promising results have been achieved, a clear and in-depth understanding of the benefit of the news sentiment to this task, as well as a comprehensive assessment of different architecture types in this context, is still lacking. Herein, we conduct an evaluation study that compares 3 different LLMs, namely, DeBERTa, RoBERTa and FinBERT, for sentiment-driven stock prediction. Our results suggest that DeBERTa outperforms the other two models with an accuracy of 75% and that an ensemble model that combines the three models can increase the accuracy to about 80%. Also, we see that sentiment news features can benefit (slightly) some stock market prediction models, i.e., LSTM-, PatchTST- and tPatchGNN-based classifiers and PatchTST- and TimesNet-based regression tasks models.
1 Introduction
Stock price prediction has been an active research topic due to its importance in investment strategies and financial risk assessment. As the financial market fluctuates significantly due to multiple economic, technical, and political factors, the precise prediction of stock prices remains extremely challenging. One research line has proposed to leverage news sentiment analysis, suggesting that sentiment-driven reactions are directly correlated with the variation of the stock price [2, 9].
The introduction of large language models (LLMs) such as BERT [3], revolutionized NLP, including text sentiment analysis, by facilitating deep contextual representation learning. Consequently, Domain-specialized variants, such as FinBERT [1] have achieved state-of-the-art results on well-known financial sentiment datasets [13] compared to traditional machine learning approaches [1][4].
Nonetheless, these pioneering works on LLM-based sentiment analysis for stock prediction are still in their infancy, necessitating a deeper and more extensive understanding of their capabilities and limitations. In fact, (1) existing works tend to investigate one individual LLM independently or without comparing it to other models on the same labeled financial dataset; (2) although LLM sentiment outputs can be encoded under different representations, such as probability scores, and discrete labels, a study including these different representations and daily aggregation methods is missing; and (3) the fusion of sentiment analysis outputs with time-series stock data is underexplored, especially with recent state-of-the-art architectures such as patch-based transformers [8], temporal graph neural networks [6], and temporal convolution models [12].
To fill these gaps, this paper presents an evaluation study that assesses the use of different LLM-based sentiment analysis approaches for stock prediction. In particular, it compares three LLM architectures, including FinBERT [1], RoBERTa [7], and DeBERTa [4], for financial sentiment classification. The latter are then integrated with traditional and recent state-of-the-art time-series models for stock prediction, such as Long Short-Term Memory (LSTM) networks [5], TimesNet [12], PatchTST [8], and tPatchGNN [6]. The obtained results on well-known datasets demonstrate that the sentiment analysis models are complementary and can achieve about 80% of accuracy when combined into an ensemble model, i.e. by training an SVM model on their outputs. The obtained news sentiment features can improve the stock price predictions, particularly in classification tasks, improving the accuracy of LSTM-, tPatchGNN- and PatchTST-based classifiers. It also improves the performance of regression models based on TimesNet and PatchTST architectures.
2 Proposed Experimental Study
Sentiment analysis models:
We have included 6 sentiment analysis models: 3 well-established Transformers-based ones and three ensemble models that we trained using the combination output of the three former. The transformers-based models are: FinBERT [1], a domain-adapted variant of BERT [3] finetuned on large-scale financial corpora; RoBERTa [7] and DeBERTa [4], two general-purpose optimised versions of BERT, which perform efficiently on finance-specific tasks [11]. For a given input text, they output a sentiment class (negative, neutral or positive) and the prediction confidence score. They are deterministic by default, non-generative and thus less prone to hallucinations, enforcing the reproducibility of our work and reducing the impact of randomness on our results. Using their predictions on a subset of SEntFiN [10] dataset, We trained three ensemble models, via traditional machine learning algorithms: Random Forest (RF), Logistic Regression (LR), and Support Vector Machines (SVM).
Stock prediction methods:
For the downstream task of stock movement prediction, we adopt four time-series forecasting methods representing fundamentally different architecture families. Long Short-Term Memory (LSTM) network [5] is included as a baseline for sequential modeling. PatchTST [8] and TimesNet [12] represent transformer-based architectures optimized for multivariate time-series analysis, with PatchTST utilizing a patching mechanism for temporal segmentation and TimesNet employing temporal variation modeling in two-dimensional space. tPatchGNN [6] extends these capabilities through the integration of temporal patching with graph neural networks, enabling the capture of both temporal dynamics and relational dependencies between assets.
Datasets:
The study examines five major publicly traded companies – Microsoft (MSFT), Amazon (AMZN), Apple (AAPL), Netflix (NFLX), and Tesla (TSLA) – over the period from March $10^{th}$ 2022 to April $2^{nd}$ 2025. We collect the related daily stock data from Yahoo Finance and the corresponding news (total of over 96000) via the AlphaVantage API. News are assigned to trading days according to their release timestamps, with weekend and holiday publications mapped to the next available trading day. Any missing values in the price series are handled through forward filling for continuity in time-series modeling.
In order to evaluate the efficiency, complementarity of the Financial sentiment analysis models and to train the ensemble models, we have used a labelled financial news dataset; SEntFiN 1.0 [10]. Unlike the Alphavantage dataset, SEntFiN provides news as text-sentiment pairs (over 10,700 news headlines), enabling model evaluation, but without any information about their issuing date, or related stock, which makes it unsuitable for the rest of the study.
Proposed daily sentiment representations:
Because multiple news items may be released on the same trading day, we apply aggregation procedures to group individual article-level sentiments into a single daily feature set. To do so, we compute the daily sum of sentiment scores to capture overall sentiment intensity, the minimum and maximum scores to reflect the most pessimistic and most optimistic signals of the day, and the majority-vote class to summarise the prevailing sentiment. The resulting dataset contains both market-based variables (daily volume and closing price) and sentiment-based variables in each aggregated representation format.
The prediction targets
are defined in four forms: price, factor, delta, and binary trend movement. Formally, let $P_{t}$ denote the closing price at trading day $t$, and let $\tilde{P}_{t}=P_{t+1}$ represent the next-day closing price aligned to time $t$. The factor (price ratio) is computed as $\text{Factor}_{t}=\frac{\tilde{P}_{t}}{P_{t}}=1+r_{t}.$, the delta (price change) as $\Delta_{t}=\tilde{P}_{t}-P_{t}=P_{t+1}-P_{t}.$ and the Binary (upward vs. downward movement) as $\text{Binary}_{t}=\mathbb{1}[\tilde{P}_{t}>P_{t}]=\begin{cases}1,&\tilde{P}_{t}>P_{t},\\ 0,&\tilde{P}_{t}\leq P_{t}.\end{cases}$
Metrics:
Predictive performance is quantified using classification and regression metrics, depending on the target formulation. For instance, to evaluate the efficiency in predicting directional (up or down) stock trends, we compute the Accuracy, F1-score, and the Area Under the ROC Curve (AUC). For regression tasks, i.e. predicting the stock price change, we compute the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and the coefficient of determination $R^{2}$ to approximate the variance proportion explained by the model.
Implementation details:
We use the LLM-based sentiment analysis model with their default recommended setup. In order to set a fair base of comparison and mitigate data-leakage threats, we exclude news from our dataset that has been used in training or fine-tuning these models. Ensemble model classifiers are trained on 80% of SentFin dataset (80-20 Training-test split).
The stock market datasets are temporally split into subsets of 70% for training, 10% for validation, and 20% for test to avoid look-ahead bias and ensure that model evaluation reflects real-world forecasting conditions. The data is structured into fixed-length sequences (rolling window) of 30 consecutive days for training, enabling the models to capture temporal dependencies over approximately one month of trading activity. Prior to training, all numerical features are normalised (Min-Max). Distinct stock prediction models have been trained (10 times each with different random seeds), offering all possible combinations of considered forecasting architectures for every target stock, for every prediction task, using or not using sentiment features from every sentiment analysis model.
We conduct the experiments on a standard computing environment [^1] NVIDIA GeForce GTX 1650 Ti GPU with 4 GB VRAM, a 2.60 GHz CPU and 16 GB of RAM. The system is equipped with CUDA 12.7 and driver version 565.77.01.. We implement the experimental study pipelines in Python using well-known open-source data manipulation and machine learning libraries. We make our complete codebase and setup instructions with additional details available  [^2] https://github.com/Walids35/capstone-stock-prediction to enable reproducibility and support similar future research.
3 Results and discussion
3.1 Effectiveness and complementarity of sentiment analysis models
Figure 1: Venn diagrams showing overlap among Original labels, FinBERT, RoBERTa, and DeBERTa across sentiment classes.
Although being a generic-purpose LLM, DeBERTa outperforms both RoBERTa and FinBERT, predicting sentiments accurately for over 75% of the dataset, followed by FinBERT then RoBERTa scoring both below 70%; 70.9% and 58.9%, respectively. The overlap and disjoint sets of accurately predicted news presented in the Venn diagrams of Figure 1 endorse the complementarity between the models, showing that they do not accurately predict the same news of the same classes. This motivates the training of a learning approach which would learn how to combine results from each model in order to improve the sentiment prediction performance. The results reported in Table 1 endorse this idea, showing an advantage in employing an ensemble model over using any LLM separately, i.e. an SVM ensemble model achieves on average about 80% of accuracy, precision, recall and F1 score.
Table 1: Performance of ensemble classifiers trained on combined outputs from FinBERT, RoBERTa, and DeBERTa.
Model Accuracy Precision Recall F1-Score
FinBERT 0.696 0.713 0.701 0.7
RoBERTa 0.589 0.744 0.591 0.585
DeBERTa 0.752 0.761 0.759 0.755
Logistic Regression 0.778 0.791 0.778 0.782
Random Forest 0.788 0.796 0.788 0.790
SVM 0.791 0.799 0.791 0.793
3.2 Impact of news sentiment on stock movement prediction
[…]
LSTM WSavg 0.5747 / 0.5557 6.1331 / 1.4726 0.0334 / 2.1242 15.8240 / 0.7455
LSTM NS 0.5694 / 0.5453 6.0629 / 1.4615 0.0324 / 2.0703 12.0457 / 0.5769
LSTM WSsvm 0.5720 / 0.5530 6.0971 / 1.4642 0.0330 / 2.1033 14.2921 / 0.6817
PatchTST WSavg 0.5525 / 0.4997 0.2638 / 2.9300 0.2227 / 3.1490 0.3768 / 3.6745
PatchTST NS 0.5558 / 0.5105 0.3474 / 3.5419 0.3906 / 5.0298 0.2419 / 2.3809
PatchTST WSsvm 0.5593 / 0.5137 0.3056 / 3.2822 0.2052 / 2.8807 0.3400 / 3.3225
TimesNet WSavg 0.5656 / 0.5191 0.2609 / 2.8192 0.2934 / 3.9363 0.2932 / 3.0241
TimesNet NS 0.5683 / 0.5209 0.4398 / 4.5547 0.4733 / 6.2774 0.1767 / 1.9371
TimesNet WSsvm 0.5657 / 0.5099 0.2663 / 2.7926 0.3063 / 4.0162 0.2648 / 2.7362
tPatchGNN WSavg 0.5242 / 0.4914 0.1423 / 1.5588 0.1771 / 2.2583 0.2734 / 2.7610
tPatchGNN NS 0.5165 / 0.4664 0.1367 / 1.5147 0.1703 / 2.1881 0.1971 / 2.0581
tPatchGNN WSsvm 0.5117 / 0.4866 0.1426 / 1.5617 0.1782 / 2.2708 0.2559 / 2.5903
In Table 2 we report the average stock movement prediction results of the studied architectures, achieved without news sentiment features (NS), using the SVM ensemble model predicted sentiments (WSsvm) and the average performance when using sentiment features of the 6 considered sentiment analysis classifiers (WSavg). Overall, the results show that the sentiment news has very low impact on the obtained results. Moreover, we notice that some architectures are better fitted for classification (up or down trend prediction), and others for regression (price prediction). For instance, LSTM predicts the movement trends the most accurately (57.47%), while it performs the worst in the regression tasks, yielding the highest MAEs. Overall, we observe that the sentiment features improve (slightly) the stock movement predictions of models trained on the studied architectures, i.e. PatchTST and TimesNet in regression tasks and all classifiers except TimesNet-based ones.
4 Conclusion
In this work, we investigate the financial news sentiment analysis performance in practice, particularly its impact on stock price movement prediction. We train and compare the performance of SOTA fundamentally different stock price movement prediction architectures, with different sentiment analysis models and different daily news sentiment representations. The results show that the studied sentiment analysis models perform efficiently, with DeBERTa outperforming both FinBERT and RoBERTa. Moreover, combining the three models, i.e. via an SVM algorithm, offers more accurate predictions (about 80%).When introduced in the stock price prediction tasks, the sentiment information improved the regression performance of models based on PatchTST and TimesNet, as well as the accuracy and AUC of LSTM-, PatchTST and tPatchGNN-based classifiers.
Acknowledgment
This work is supported by the Luxembourg National Research Funds (FNR) through the FinInclusion project grant NCER22/IS/16570468/NCER-FT.